# Quantifying Geospatial in the Common Crawl Corpus

### v1

@@ 4. Findings
Our estimate of the prevalence of geospatial data in CC documents is shown in Figure 2. Overall, when analysing the three CC releases from between 2019 and 2024, we find that 18.3% of the documents contain either a street address or a pair of coordinates (or both), 15.6% of the documents contain a street address, 7.1% of the document contain a pair of coordinates, and 4.4% of the documents have both an address and a pair of coordinates. These estimates are fairly consistent between the releases.
Almost 28% of the web documents that contain an address also contain a pair of coordinates, and almost 62% of the documents that contain coordinates also contain an address. (It is worth noting that the coordinate and the address found in the same document do not necessarily refer to the same location.)
The most pronounced single source of coordinates in web documents is Google Maps URLs that encode latitude and longitude (occasionally with the address, too). Examples of these can be found in the Identifying coordinates and addresses subsection.
On several occasions, Gemini judged that an IP address or a phone number or a named landmark (such as a theatre) is sufficient to locate the place. In some cases, it also returned the astronomical coordinates of celestial bodies. However, these did not meet our definition of geospatial described in the What constitutes geospatial subsection, so we excluded them from the final count.
@@ 5. Conclusion
In this paper, we analysed the contents of CC for the presence of geospatial data. Using Gemini and manual verification, we estimated that about 18.3% of all documents in recent CC releases contain a street address or a pair of coordinates. Thus, we conclude that geospatial data is fairly well represented in mostly unstructured CC documents.
Therefore, it is not surprising that LLMs exhibit some geospatial capabilities, such as the knowledge of distances and directions between landmarks and other locations. We suggest further research into the accuracy and quality of the geospatial data in CC documents is needed to better understand the geospatial biases that LLMs exhibit. We also recognise that coordinates and addresses, being specific point-level bits of information, are not the only way LLMs learn the physical space; the context of the web pages where geospatial data is found is also very important. This may include textual descriptions of places, both objective and subjective, such as those describing the surroundings or the experience of the visitors or the activities happening in the area.
Given the scale of CC, which advertises itself as having the total of ‘over 250 billion documents’232323https://commoncrawl.org/, we can estimate that 45 billion pages are likely to contain geospatial data. In some cases, the geospatial data will be context rich, making CC a potential goldmine for domain-specific geospatial datasets that can be used to train or fine-tune LLMs on geospatial tasks.
@@ Acknowledgements.
- Baack (2024) Stefan Baack. 2024. Training Data for the Price of a Sandwich: Common Crawl’s Impact on Generative AI. (Feb. 2024).
@@
- Crawl (2024) Common Crawl. 2024. Common Crawl - Open Repository of Web Crawl Data. https://commoncrawl.org/
@@
- Ilyankou et al. (2024) Ilya Ilyankou, James Haworth, and Stefano Cavazzi. 2024. CC-GPX: Extracting High-Quality Annotated Geospatial Data from Common Crawl. http://arxiv.org/abs/2405.11039 arXiv:2405.11039 [cs].

### v2

@@ 4. Findings
Our estimate of the prevalence of geospatial data in CC documents is shown in Figure 3. Overall, when analysing the three CC releases from between 2019 and 2024, we find that 18.7% of the documents contain either a street address or a pair of coordinates (or both), 16.1% of the documents contain a street address, 7.0% of the document contain a pair of coordinates, and 4.3% of the documents have both an address and a pair of coordinates. These estimates are fairly consistent between the releases.
Around 27.1% of the web documents that contain an address also contain a pair of coordinates, and 62.0% of the documents that contain coordinates also contain an address. (It is worth noting that the coordinate and the address found in the same document do not necessarily refer to the same location.)
Geospatial data occurrence by language
Figure 4 demonstrates that the frequency of geospatial data occurrence is fairly consistent between documents in English and documents in other languages. While English-language documents are slightly more likely to contain coordinates than non-English-language documents (7.5% vs 6.0%), the latter are slightly more likely to contain addresses (16.2% vs 15.1%).
The most pronounced single source of coordinates in sampled web documents is URLs to Google Maps, present either as a hyperlink to the mapping service itself (<a href="...">), or as an embedded map (<iframe src="...">). An example of such URL can be found in Subsection 3.3. Thus, the fact that Google Maps is not widely used in some major non-English speaking countries such as China (where Baidu Maps242424https://map.baidu.com/ and AutoNavi252525https://mobile.amap.com/ are popular), Russia (Yandex Maps262626https://yandex.ru/maps), Iran or Vietnam, is likely to have a negative impact on the frequency of coordinates in non-English-language web documents.
On several occasions, Gemini judged that an IP address or a phone number or a named landmark (such as a particular theatre) is sufficient to locate the place. In some cases, it also returned the astronomical coordinates of celestial bodies. However, these did not meet our definition of geospatial described in Subsection 3.2, so we excluded them from the final count.
@@ 5. Conclusion
In this paper, we analysed the contents of CC for the presence of geospatial data. Using Gemini 1.5 Flash and manual verification on a sample of 119k documents, we estimated with 95% confidence that of web documents in recent CC releases contain a street address or a pair of coordinates. Thus, we conclude that geospatial data is fairly well represented in mostly unstructured CC documents.
Therefore, it is not surprising that LLMs exhibit some geospatial capabilities, such as the knowledge of distances and directions between landmarks and other locations. We showed that documents in the English language and from the .com TLD are disproportionately present in CC; this imbalance risks amplifying some worldviews while silencing others, exposing the language models to geospatial biases during pre-training. Therefore, we suggest the need for further research into the accuracy and quality of the geospatial data for different regions and languages, which can help us better understand the geospatial biases that LLMs exhibit.
We also recognise that coordinates and addresses, being specific point-level bits of information, are not the only way LLMs learn the physical space; the context of the web pages where geospatial data is found is also very important. This may include textual descriptions of places, both objective and subjective, such as those describing the surroundings or the experience of the visitors or the activities happening in the area.
With roughly documents potentially containing geospatial data, Common Crawl a potential goldmine for domain-specific datasets that can be used to train, fine-tune, or benchmark LLMs on geospatial tasks.
@@ Acknowledgements.
- Baack (2024a) Stefan Baack. 2024a. A Critical Analysis of the Largest Source for Generative AI Training Data: Common Crawl. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. ACM, Rio de Janeiro Brazil, 2199–2208. https://doi.org/10.1145/3630106.3659033
- Baack (2024b) Stefan Baack. 2024b. Training Data for the Price of a Sandwich: Common Crawl’s Impact on Generative AI. (Feb. 2024).
@@
- Crawl (2024a) Common Crawl. 2024a. Common Crawl - Open Repository of Web Crawl Data. https://commoncrawl.org/
- Crawl (2024b) Common Crawl. 2024b. Statistics of Common Crawl Monthly Archives by commoncrawl. https://commoncrawl.github.io/cc-crawl-statistics/plots/languages
@@
- Ilyankou et al. (2024) Ilya Ilyankou, Meihui Wang, James Haworth, and Stefano Cavazzi. 2024. CC-GPX: Extracting High-Quality Annotated Geospatial Data from Common Crawl. https://doi.org/10.48550/arXiv.2405.11039 arXiv:2405.11039 [cs].
- ISO (2012) ISO. 2012. geographic data — MLGT: The authoritative multi-lingual geographic information terminology database. https://isotc211.geolexica.org/concepts/202/
@@
- Paolillo et al. (2005) John Paolillo, Daniel Pimienta, and Daniel Prado. 2005. Measuring Linguistic Diversity on the Internet. (2005).
@@
- W3Techs (2024) W3Techs. 2024. Usage Statistics and Market Share of Top Level Domains for Websites, August 2024. https://w3techs.com/technologies/overview/top_level_domain